Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 15/8/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts : Colin/ChrisR stands
  • DC update
  • Status of alarm clearance (DC/Ajit)
  • Hydrogen system update and plans (KLong/CWhyte)
  • Reanimating the SS magnets (it's alive! Or it it?) (Alan)
  • Cycle 2017 Issues
  • Hall/Operations plan : 17-08-14-CW.xlsx
  • Proposed Ops startup schedule
    • Initial discussion on system readiness : Wednesday August 23rd
    • Pre-startup meeting : Monday Sep 4th
    • Pulser run : Sep 7th/8th
    • Beam tuning and beam test run : sometime in the week starting Sep 11th
  • AOB

Minutes

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
    • LH2: ATEX work in progress. exVeritas here Mon./Tue. Delay as we need a replacement emergency light, and we cannot remove the heaters from the shed; this won't finish by tomorrow as expected.
      The quench line is being installed; should finish tonight.
      JB: Richard Hale has completed maintenance of the external fans.
      The pipework outside the Hall has been labelled.
      The crash barrier fix is in progress.
      The Hall has been partly tidied prior to LH2 sign-off tour; another round tomorrow.
    • The Tracker cryostat refurb has completed - all are now colder than before. Calibration is in progress.
    • Sandor and Alan checked out the QPS for the SS's. SSU E-C-E-M1-M2 & SSD E-C-E-M2 have been powered to 50 A and fast ramp-down tested. Found that SSU PSU doesn't re-set max. voltage after fast ramp-down: not clear why as code is same for both. Need to test.
      On Saturday took SSU and SSD to operating current (208 A solenoid mode) and held for 30 mins.
      Limited cooling headroom on SSD. Have replacement adsorbers en route from USA, need to replace in compressors and re-gas; will take a couple of days. Failed SSD CC1 will be left where it is, spare units are available.
    • The Diffuser was tested during the SS ramp - operated OK.
    • A/C unit 3 iced up again, and has now been thawed out. Paul and Henry set up a syphon to drain the 100+ litres of water caught in a waste bin into the Trench gutter (the rest went on Hall floor). The sensor dongle detected a Trench Water leak but this failed to come up in the Alarm Handler.
      This unit was only re-commissioned last week - get experts to look at it.
  • Status of alarm clearance (DC/Ajit)
    Need to re-enable Trackers.
    Not sure what to do about VME System Temperatures.
    DS again has the falling shield pressure issue. Mark Tucker is aware, has not yet found a way to correct it.
  • Hydrogen system update and plans (JBoehm)
    The Beacon is stuck on red and alarming; Phil to add acknowledge feature to software.
    ISIS have OK'd the quench line flap valves and helium flush in principle, but needs a new control sequence.
    Leak testing is largely complete.
    We have shown we can cool and hold level liquid level with Neon.
    Post-HAZOP documentation is almost complete, needs final ATEX records.
    Meeting with Fire Officer who approved changes but wants documentation.
    Phil Warburton down next week for software changes; plan was to then run Hydrogen but now looks unlikely. Should we prepare another Neon test? No, just Helium.
    Need an LH2 documentation meeting tomorrow; already have an early Thursday LH2 meeting planned. Aggregate into one 10am meeting? Decide elsewhere.
  • Hall/Operations plan : 17-08-14-CW.xlsx
    Sandor would like to power the entire channel (inc. FC) - would need FC expert (John Cobb?). Would have to be Friday.
    It's not clear how useful this would be, compared to looking at the fast ramp-down re-set issue. JBoehm would prefer to test latter with magnets, not shorts.
    Also FC has a vacuum issue - maybe a leak somewhere? JBoehm to remove the suspect flange.
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    Yes!
  • AOB
    • JLanglands: changed a carrier board in the R78 VME crate, and the target is now running again. JL to calibrate the new laser. After the target there has been validated, the big laser and target1ctl will come back to the MLCR
    • MUchida: the Tracker volume bubblers are bubbling ... but there's no gas supply connected!

Updated by Nebrensky, Henry over 6 years ago · 8 revisions